OpenStudy (anonymous):

write the equation in the form f(x)= k^2

OpenStudy (anonymous):

Isnt that an equation? I am confused what you are asking for.

OpenStudy (anonymous):

k(3)=54 and n=2

OpenStudy (anonymous):

help

OpenStudy (agent0smith):

write the equation in the form f(x)= k^2 k(3)=54 and n=2 ^^^ You might have to rewrite the question... it's unclear what you're asking for.

OpenStudy (anonymous):

f(x)= kn^2

OpenStudy (unklerhaukus):

f(x)=kx^n ?